OpenStudy (anonymous):

In a coordinate system, points A(-2, 3) and B(6, -1) have a midpoint of X. What is the midpoint between A and X? Show your work.

OpenStudy (anonymous):

find first the coordinates of X then find the coordinates of the midpoint of AX

OpenStudy (anonymous):

confused on how to do it

hartnn (hartnn):

the formula to get the co-ordinates of midpoint(m,n) is m=(x1+x2)/2,n=(y1+y2)/2 where (x1,y1)(x2,y2)are endpoints so u can get the co-ordinates of X with A and B as endpoints now again apply this formula with A and X as endpoints....

OpenStudy (anonymous):

Just.. (-2+6) divided by 2 for ur x coordinate and.. (3+1) divided by 2 for ur y coordinate

OpenStudy (anonymous):

so its (4,2)

OpenStudy (anonymous):

yeapz

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem~